OpenStudy (hmarie29):

solve the system using an algebraic method x - y = 5 2x - 4y + 6 = 0

OpenStudy (otherworldly):

are those two together

OpenStudy (hmarie29):

yes

OpenStudy (ben00):

So without x and y it would be 2 - 4 + 6

OpenStudy (hmarie29):

answer choices r ( 2,-3) (3,2) infinite number of points or no solution

OpenStudy (anglina4196):

First you multiply the top equation by -2 to cancel out the x variable

OpenStudy (anglina4196):

so.... what is it then

OpenStudy (hmarie29):

umm (2, -3) ?

OpenStudy (anglina4196):

no multiply the top equation by -2 then it would be -2x+2y=-10 2x-4y+6=0

OpenStudy (anglina4196):

Then you would add the equations together so.... -2y+6=-10

OpenStudy (ben00):

This is interesting.

OpenStudy (hmarie29):

A. ( 2,-3) B. (3,2) C. infinite number of points D. no solution these r my only choices

OpenStudy (anglina4196):

Then solve the equation -2y+6=10

OpenStudy (hmarie29):

-2

OpenStudy (anglina4196):

i mean -2y+6=-10

OpenStudy (hmarie29):

8